### v1

Abstract.
The complexity of the Generally Accepted Accounting Principles (GAAP) and the hierarchical structure of eXtensible Business Reporting Language (XBRL) filings make financial auditing increasingly difficult to automate and verify. While large language models (LLMs) have demonstrated strong capabilities in unstructured text understanding, their ability to reason over structured, interdependent, and taxonomy-driven financial documents remains largely unexplored. To fill this gap, we introduce 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first taxonomy-aligned, structure-aware, multi-document benchmark for evaluating LLMs on financial auditing tasks. Built from real US-GAAP-compliant XBRL filings, 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: defines three complementary subtasks, FinSM for semantic consistency, FinRE for relational consistency, and FinMR for numerical consistency, each targeting a distinct aspect of structured auditing reasoning. We further propose a unified evaluation framework integrating retrieval, classification, and reasoning metrics across these subtasks. Extensive zero-shot experiments on 13 state-of-the-art LLMs reveal that current models perform inconsistently across semantic, relational, and mathematical dimensions, with accuracy drops of up to 60–90% when reasoning over hierarchical multi-document structures. Our findings expose the systematic limitations of modern LLMs in taxonomy-grounded financial reasoning and establish 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: as a foundation for developing trustworthy, structure-aware, and regulation-aligned financial intelligence systems. The benchmark dataset is available at Hugging Face [^1] FinAuditing.
XBRL auditing, Benchmark, Large language model, Information retrieval, Information extraction, Question answering
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Applied computing Extensible Markup Language (XML)$c$s: Information systems Test collectionsccs: Information systems Information extractionccs: Information systems Question answering
Figure 1. Trends in the proportion of financial restatements among US companies from 2014 to 2024, categorized by reissuance and revision restatements. Data adapted from the Financial Times report, ”Accounting errors force US companies to pull statements in record numbers” (Dec 9, 2024).
1. Introduction
The Generally Accepted Accounting Principles (GAAP) provide the foundation for corporate financial reporting in the United States, but their growing complexity and continuous updates have made compliance verification increasingly difficult (Debreceny et al., 2010; Hoitash and Hoitash, 2018). Financial restatements caused by misapplied GAAP rules have become a persistent source of reputational and regulatory risk, with recent reports indicating that severe reissuance restatements now account for the majority of filing corrections as shown in Figure. 1. While the introduction of the eXtensible Business Reporting Language (XBRL) has standardized digital disclosures, its hierarchical and interdependent structure introduces new challenges for automated auditing and anomaly detection.
With the rapid advancement of large language models (LLMs), researchers have increasingly explored their applicability across a wide range of tasks in the financial domain, including information extraction (IE) (Loukas et al., 2022; Sharma et al., 2023; Wang et al., 2025), text generation (TG) (Mukherjee et al., 2022; Xie et al., 2023), text analysis (TA) (Sinha and Khandait, 2021), question answering (QA) (Chen et al., 2021, 2022; Zhu et al., 2021; Zhao et al., 2022, 2023), risk management (RM) (Feng et al., 2023), and forecasting (FO) (Soun et al., 2022), showing potential to improve GAAP interpretation and understanding in XBRL filings. However, as summarized in Table 1, existing benchmarks mainly rely on unstructured or semi-structured data drawn from a single document or limited contexts, such as financial reports and news, while overlooking inter-dependencies, compositional hierarchies, and relational constraints that characterize real-world XBRL documents. Moreover, most financial benchmarks assess LLMs on isolated understanding and reasoning abilities, focusing on surface-level semantic understanding within plain texts or single documents. They rarely evaluate deeper capabilities such as structured semantic retrieval, hierarchical relation understanding, and multi-step reasoning, which are fundamental for capturing the structural and logical complexity inherent in financial data. Furthermore, while benchmarks such as FinTagging (Wang et al., 2025) initially utilize financial taxonomies, most do not fully leverage their hierarchical and relational structures, leaving taxonomy-driven reasoning largely unexplored. This omission limits the ability of current evaluations to measure how well LLMs can reason over structured, taxonomy-driven, and cross-document dependencies, which are crucial for a comprehensive understanding of finance.
Table 1. Comparison of existing financial NLP benchmarks and 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:. The table shows whether each benchmark supports structured retrieval (Struct.Retrieval), hierarchical information extraction (Hiera.IE), multi-step numerical reasoning (MultiStep.Reasoning), taxonomy-driven supervision (Need.Taxonomy), and at the cross-document level (At.Cross-Doc).
Benchmark Source Data Format Task Struct.Retrieval? Hiera.IE? MultiStep.Reasoning? Need.Taxonomy? At.Cross-Doc?
FiNER (Loukas et al., 2022) Financial reports Text IE
FNXL (Sharma et al., 2023) Financial reports Text IE
Headlines (Sinha and Khandait, 2021) Financial news Text TA
ECTSum (Mukherjee et al., 2022) Earnings call transcripts Text TG
CCF (Feng et al., 2023) Credit card transactions Text RM
BIGDATA22 (Soun et al., 2022) Tweets, historical prices Text FO
FinTagging (Wang et al., 2025) Financial reports Text & Table IE, IR
FinQA (Chen et al., 2021) Financial reports Text & Table QA
ConvFinQA (Chen et al., 2022) Financial reports Text & Table QA
TAT-QA (Zhu et al., 2021) Financial reports Text & Table QA
MultiHiertt (Zhao et al., 2022) Financial reports Semi-structured and hierarchical Text&Table QA
DOCMATH-EVAL (Zhao et al., 2023) Financial reports Semi-structured document segmentation QA
FinAuditing (ours) XBRL filings Structured and hierarchical multi-document segmentation IE, IR, QA
To address these gaps, we propose 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first benchmark that evaluates LLMs on structured semantic retrieval, hierarchical relation understanding, and multi-step reasoning over taxonomy-driven and interconnected XBRL filings, focusing on detecting inconsistencies and errors across structured financial documents. To this end, 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: encompasses three independent tasks: the financial semantic matching (FinSM) task, which examines whether models can detect semantic inconsistencies between financial disclosures and standardized taxonomy concepts; the financial relationship extraction (FinRE) task, which evaluates the ability to identify structural and hierarchical errors among related financial elements; and the financial mathematical reasoning (FinMR) task, which measures whether models can reason over numerical relationships to verify the logical correctness of reported values. To comprehensively assess these tasks, we design task-specific evaluation metrics that collectively capture retrieval accuracy, structural understanding, and numerical reasoning performance, using measures such as Hit Rate, Macro-F1, and error-rate indicators for a multi-perspective evaluation of LLM capabilities. As a result, our work introduces a more comprehensive and systematic evaluation framework that measures LLMs’ capabilities on structured, hierarchical, and multiple financial documents.
We conduct an extensive evaluation on 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: under the zero-shot setting, covering 13 state-of-the-art (SOTA) proprietary and open-source LLMs across three structured auditing tasks. Our results show that while leading models such as DeepSeek-V3 in FinSM, GPT-4o in FinRE, and Fin-o1-14B in FinMR achieve the highest performance within their respective tasks, all models exhibit substantial limitations. DeepSeek-V3 demonstrates better retrieval accuracy yet struggles with coverage and semantic alignment across interconnected structured documents. GPT-4o shows strong hierarchical relation understanding across all three relation types, whereas most open-source models have difficulty with the “CombinationErr” type. Fin-o1-14B, despite its domain specialization, remains weak in schema interpretation and structural consistency during numerical reasoning. Overall, these findings highlight that even frontier models fall short of reliable structured auditing, revealing persistent challenges in semantic retrieval, relational reasoning, and multi-step computation over hierarchical XBRL data.
We conclude our main contributions as follows:
[…]
2. Related work
Existing financial NLP research has made progress in information extraction (Loukas et al., 2022; Sharma et al., 2023; Wang et al., 2025), semantic understanding (Mukherjee et al., 2022; Xie et al., 2023; Sinha and Khandait, 2021), and numerical reasoning (Chen et al., 2021, 2022; Zhu et al., 2021; Zhao et al., 2022, 2023), yet most benchmarks focus on unstructured or semi-structured data from single documents, such as financial reports or news articles, overlooking the structured and interconnected nature of XBRL filings.
Early work on XBRL mainly targeted localized tasks. Loukas et al. (Loukas et al., 2022) developed an XBRL tagging benchmark for numerical entity extraction, and Sharma et al. (Sharma et al., 2023) proposed numeral labeling within XBRL sentences. Recently, Wang et al. (Wang et al., 2025) explored structured tag normalization via LLMs based on the US-GAAP taxonomy. Steven et al. (Katz et al., 2024) employed ChatGPT for tag retrieval, while Han et al. (Han et al., 2024) introduced XBRL-Agent to assess LLMs’ analytical and numerical abilities on XBRL data. However, these studies evaluate narrow subtasks and fail to capture the hierarchical and relational dependencies that define real-world XBRL structures.
Beyond XBRL, benchmarks such as FinQA (Chen et al., 2021), ConvFinQA (Chen et al., 2022), TAT-QA (Zhu et al., 2021), MultiHiertt (Zhao et al., 2022), and DocMath-Eval (Zhao et al., 2023) assess reasoning over textual and tabular data, yet they remain text-centric and lack taxonomy-driven supervision. To fill this gap, our proposed 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: provides the first benchmark to evaluate LLMs on structured, hierarchical, and cross-document reasoning grounded in real XBRL filings.
3. 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:
Detecting errors across the hierarchical and structured financial documents requires understanding semantics, hierarchical relations, and numerical logic. As shown in Figure 2, to evaluate these capabilities, we propose 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, a benchmark that measures how well LLMs identify semantic, structural, and numerical inconsistencies across multi-document XBRL filings.
Figure 2. Overview of the 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark framework for evaluating error detection on XBRL filings across three tasks.
3.1. Preliminary knowledge
The US-GAAP Taxonomy serves as a standardized vocabulary and hierarchical framework that defines both the financial concepts (e.g., us-gaap:CashAndDueFromBanks) and the relationships among them. It provides a consistent semantic structure for financial reporting, ensuring comparability and enabling automated analysis.
An XBRL filing is a structured digital financial report formatted in XML. It consists of multiple interconnected documents, including an instance document, schema document, and several linkbases, which together define and link the representations and relationships of financial elements (see Table 2 for details).
Additionally, Fact corresponds to the reported values associated with US-GAAP tags in the instance documents. Context specifies details about a reported fact (e.g. reporting period). Axis represents a categorical dimension that allows financial value to be classified from different perspectives. For instance, ”RevenueByGeographyAxis” organizes revenue based on geographic regions. Member is a specific value within an Axis, representing a particular category. For example, within ”RevenueByGeographyAxis”, ”USMember” denotes revenue generated in the United States, while ”EuropeMember” refers to European revenue.
Table 2. Main Components of XBRL Filings
Component Description
Instance Document Actual financial data structured according to a predefined taxonomy.
Schema Document Defines the structure, classification, and data types of financial elements.
Calculation Linkbase Specifies arithmetic relationships between elements to ensure logical consistency.
Presentation Linkbase Organizes financial data into a human-readable structure for reporting.
Definition Linkbase Captures semantic relationships between elements, such as component relations.
Label Linkbase Provides human-readable names and descriptions, supporting multiple languages.
3.2. Task Formulation
Given a target XBRL filing with six interconnected, XML-format documents $\mathcal{D}=\{d_{1},d_{2},d_{3},d_{4},d_{5},d_{6}\}$ and a reference structured document (US-GAAP taxonomy) $\mathcal{T}$, where $d_{i}$ refers to one of the documents described in Table 2. We aim to enable LLMs to identify errors $\mathcal{E}$, including semantic matching errors, relationship extraction errors, and mathematical reasoning errors, in the target filing by understanding its internal structure and semantics in conjunction with domain knowledge derived from the reference taxonomy. We formulate it as follows:
(1) $f:\left(\mathcal{D},\mathcal{T}\right)\mapsto\mathcal{E}$
where $\mathcal{E}$ is the specific description of an error found in the $\mathcal{D}$.
To operationalize error detection across the key dimensions of XBRL filings, we identify three fundamental capabilities required for effective model evaluation: semantic consistency, structural relationship understanding, and numerical reasoning. These capabilities correspond to the primary sources of errors observed in real-world filings and capture the essential challenges in interpreting and validating structured financial data. Based on this decomposition, FinAudit defines three tasks: FinSM for identifying mismatches between textual mentions and standardized taxonomy labels, FinRE for detecting errors in hierarchical or compositional relationships among financial items, and FinMR for verifying the correctness of numerical values based on accounting logic and contextual inference. Each sub-task is associated with a specific input format tailored to its objective.
3.2.1. Financial Semantic Matching (FinSM)
The semantic matching task focuses on identifying financial tags in an XBRL filing that do not correctly align with standardized concepts defined in the US-GAAP taxonomy. The objective is to assess whether LLMs can accurately detect mismatched US-GAAP tags by filtering out semantically inconsistent or inappropriate concept assignments from a set of complex, structured documents.
This task follows the information retrieval paradigm: given a query $Q$ describing a financial term that represents either currency or concentration of credit risk, an XBRL filing $\mathcal{D}$, and a US-GAAP taxonomy $\mathcal{T}$. We formulate this task as follows:
(2) $f:(Q,\mathcal{D},\mathcal{T})\mapsto\mathcal{E}$
where $\mathcal{E}$ is the set of mismatched US-GAAP tags after retrieving.
FinSM Example You are an auditor for XBRL filings. Given the question and the provided filing (schema, presentation, calculation, definition, label, instance, and US-GAAP taxonomy), identify erroneous US-GAAP concepts. You must reference the US-GAAP taxonomy to retrieve and check elements across the entire filing. Output only a JSON array of strings, each string is one erroneous concept. Output Example: ["us-gaap:Revenues", "us-gaap:OperatingIncomeLoss"] No explanations. No extra text. Only the JSON array.
3.2.2. Financial Relationship Extraction (FinRE)
The relationship extraction task aims to identify the structural relationships among financial elements in a filing. It assesses an LLM’s ability to interpret hierarchical and compositional dependencies by aligning the document structure with the external taxonomy. This task reflects the LLMs’ capacity for semantic understanding and structured data interpretation within the context of financial reporting.
FinRE Example You are an auditor for XBRL filings. Given two concepts and the provided filing (schema, presentation, calculation, definition, label, instance, and US-GAAP taxonomy), determine which one of the following erroneous relationship types best describes the relationship between the two concepts: (1) Reversal Definition: <definition>. (2) Inappropriateness Definition: <definition>. (3) CombinationErr Definition: <definition>. You must reference the US-GAAP taxonomy to retrieve and check the two concepts across the entire filing. Output only one of the following labels exactly, chosen strictly from:["Reversal", "Inappropriateness", "CombinationErr"]. No explanations. No extra text. Only the label.
More specifically, given a relationship label space $L=\{l_{1},l_{2},l_{3}\}$ with the definitions for each relationship label, an XBRL filing $\mathcal{D}$, and a US-GAAP taxonomy $\mathcal{T}$. The goal of this task is to classify the relation type $r$ for two specific elements $e_{1}$ and $e_{2}$ based on $\mathcal{D}$ and $\mathcal{T}$. More specifically, we define three relation error types in $L$: (1) $l_{1}=Reversal$, occurs when the hierarchical relationship between sibling and child elements is mistakenly reversed; (2) $l_{2}=Inappropriateness$, indicates that a child element is incorrectly associated with an inappropriate parent element; and (3) $l_{3}=CombinationErr$, refers to an invalid combination of axis and member elements that violates the structural constraints defined in the taxonomy. Therefore, we formulate this task as follows.
(3) $f:(e_{1},e_{2},L,\mathcal{D},\mathcal{T})\mapsto\mathcal{E}$
where $\mathcal{E}$ is the specific identified relation type $r$ from the relationship label space $L$.
3.2.3. Financial Mathematical Reasoning (FinMR)
The mathematical reasoning task focuses on inferring numerical logic among financial elements based on structured representations, calculation hierarchies, and logical constraints defined in the XBRL filing. The objective is to assess whether reported values are internally consistent and compliant with domain-specific standards.
Specifically, given two questions $q_{1}$ and $q_{2}$, where $q_{1}$ concerns the extraction of a reported value and $q_{2}$ pertains to the calculation of the corresponding real value, an XBRL filing $\mathcal{D}$, and a US-GAAP taxonomy $\mathcal{T}$, the task is to extract the reported value $v$ for a given instance in $\mathcal{D}$ and to compute the numeric value $\mu$ for that instance, which is then used to verify whether the reported value $v$ is correct.
(4) $f:(q_{1},q_{2},\mathcal{D},\mathcal{T})\mapsto\mathcal{E}$
where $\mathcal{E}$ denotes the generated answer containing the extracted reported value $v$ and the computed value $\mu$, represented in JSON format.
FinMR Example You are an auditor for XBRL filings. Given the question and the provided filing (schema, presentation, calculation, definition, label, instance, and US-GAAP taxonomy), identify the reported value of a financial element and calculate the actual value that should be reported based on calculation relationships. Answer strictly in the following JSON format: { "extracted_value": "<numeric value reported in the instance document, or 0 if not found, keep the same format as in the XBRL filing>", "calculated_value": "<numeric value computed from calculation relationships, or 0 if not computable, use the same number formatting style as extracted_value>" } No explanations. No extra text. Only the JSON object.
4. Data Annotation and Evaluation Setting
4.1. Raw Data Collection
To create this benchmark, we collected a total of 372 XBRL filings, which were processed through two steps: pre-screening and filtering. Pre-screening serves as the foundation for filtering, enabling the identification and collection of common errors in XBRL filings.
For the pre-screening step, we accessed the XBRL US official website [^2] https://xbrl.us/data-quality/filing-results/ and randomly collected 4,545 error messages from 372 companies spanning the years 2020 to 2024. These messages, linked to the DQC US rules [^3] https://github.com/DataQualityCommittee/dqc_us_rules/tree/master, were analyzed to assess rule coverage and to identify the most common error types. Our analysis revealed that the 4,545 messages contain more than 100 distinct error types. However, as shown in Table 3, 9 error types occur most frequently, accounting for approximately 60.33% of all cases. These errors can be grouped into three broad categories: semantic-based, relation-based, and calculation-based.
For the filtering step, we extracted 1,102 error messages from the initial set of 4,545 as the first version of our benchmark, restricting the scope to data from 2022 to 2024 and selecting only those associated with the 9 common error types listed in Table 3 with corresponding DQC US IDs. Leveraging the “sec-url” field in these reports, we then accessed the SEC (U.S. Securities and Exchange Commission) [^4] https://www.sec.gov/ database to obtain 218 corresponding XBRL filings, which constitute the foundation of our benchmark raw data.
Table 3. Distribution of common error types in XBRL filings.
Broad Category Error Type # Cases Proportion (%) DQC US ID
Semantic-based FS with no associated calculation 386 8.49 DQC_0099
Concentration risk 217 4.77 DQC_0109
Location axis with a single member 208 4.58 DQC_0137
Relation-based Sibling and child relationships 570 12.54 DQC_0081
Axis with inappropriate members 313 6.89 DQC_0001
Inappropriate cash flow presentation 219 4.82 DQC_0145
Calculation-based FS calculation check with no dimensional data 260 5.72 DQC_0126
Negative values 368 8.10 DQC_0015
FS tables dimensional cross check 202 4.44 DQC_0117
Total 2,743 60.33 -
4.2. Data Annotation
Based on the filtered raw data described in Section 4.1, we annotated a total of 1,102 instances for the 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark. As the first benchmark to focus on structured, multi-document, and taxonomy-driven inputs, 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: provides a unique setting for evaluating LLMs. Table 4 shows that each task involves long and complex inputs, with average lengths exceeding 33k tokens per instance when measured using the “cl100k_base” GPT tokenizer. Among them, FinRE contains the largest total token volume and number of instances, while FinMR exhibits the longest average input length per instance. These statistics highlight the structural and relational complexity of XBRL filings, underscoring the need for careful annotation to ensure reliable evaluation.
Table 4. Statistics of input length (tokens) across tasks. Tokens are calculated using the “cl100k_base” tokenizer.
Task Total Tokens Max Tokens Avg. Tokens Std. Tokens Instances
FinSM 11,170,024 62,777 33,848.56 11,402.49 330
FinRE 14,819,627 65,030 33,680.97 10,847.23 440
FinMR 11,845,255 59,641 35,678.48 9,852.73 332
Total 37,834,906 65,030 34,338.64 10,700.82 1,102
4.2.1. Data Segmentation
For each complete XBRL filing, we extracted all statement IDs, which are their associated roleType IDs from the schema document, to segment the filing into distinct sections. Each section corresponds to a specific financial statement, such as the balance sheet or cash flow statement. Using these statement IDs, we then extracted the related content across linkbase documents. In particular, the presentation linkbase defines the logical and visual arrangement of financial elements within each statement and connects the facts in the instance document to their structured representation in reports. Following this process: from the schema document to the presentation linkbase, and subsequently to other linkbases and instance documents, we segmented each filing into multiple sections. These sections, treated as sub-filings of the original XBRL filing, preserve the inherent structure while reducing document length, thereby improving data manageability and enhancing the efficiency of LLM processing and evaluation.
4.2.2. Annotation with LLM Judgment
Each XBRL filing is retrieved using the ”sec-url” from the error messages, which provide explicit details, including the specific document components (e.g., calculation linkbase or instance document) where errors occur, the incorrect values, the invalid relationships, and the expected correct results. Therefore, we do the following processing for each task using the GPT-4o-mini model. We provide the specific prompts for each task in Appendix A.
(1) For the FinSM task, we extract all reported incorrect concepts from the error messages as the ground-truth answers and retrieve the corresponding sub-filing segments based on these concepts to serve as the input XBRL filing.
(2) For the FinRE task, we map the three defined relationship types to the corresponding error messages based on the DQC US ID. Specifically, DQC_0081 corresponds to “Reversal”, DQC_0001 to “CombinationErr”, and DQC_0145 to “Inappropriateness”. For each relationship type, we extract the participating elements from the error messages and construct relationship triples in the format: (head element, relation, tail element). The head and tail elements are used as inputs, while the relation serves as the ground-truth label. Similarly, based on the extracted elements, we retrieve the corresponding sub-filing segments to serve as the input XBRL filing.
(3) For the FinMR task, based on our task formulation, we annotate two key aspects for each question: the reported value in the instance document and the inferred generative value. Both answers are derived from the error message. Since the question in this task is concept-dependent, we extract the relevant concept and period from the error message to construct the input question. Similarly, we retrieve the corresponding sub-filing segments based on the extracted concepts to serve as the input XBRL filing.
4.2.3. US-GAAP Taxonomy Chunk
To facilitate structured access to the US-GAAP Taxonomy, we divided the taxonomy into concept-centric chunks. Each concept and its associated semantic information were extracted from the official taxonomy Excel files (e.g., GAAP Taxonomy 2024.xlsx). Specifically, we parsed all relevant sheets, including Concepts, Presentation, Calculation, Definition, Reference, and Enumeration, and organized them into two major chunk types: core chunks, which contain each concept’s intrinsic attributes such as label, type, balance, period type, and documentation, and relation chunks, which describe its structural and semantic links to other concepts. The relation chunks encode various dependency types, including hierarchical presentation paths, parent–child calculation relationships, definitional arcs, and external references. The details are shown in Appendix B. This chunking process preserves the hierarchical and relational structure of the taxonomy while enabling fine-grained indexing, retrieval, and reasoning at the concept level.
4.2.4. Quality Control
To ensure the reliability and consistency of the 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark, we conducted a series of quality control procedures across both data extraction and annotation phases. (1) A random sample of 50 filings was manually checked to verify that the extracted statement segments matched their roleType identifiers and preserved the original hierarchy. (2) For annotation, 50% of the instances were double-validated by comparing GPT-4o-mini outputs with independent human review, and discrepancies were corrected to ensure accuracy. (3) Finally, taxonomy chunks were cross-checked to ensure that every relation chunk referenced a valid core concept. These steps collectively guarantee structural fidelity and reliable ground-truth annotations for evaluating LLM reasoning over structured financial data.
4.3. Evaluation Metrics
Our goal is to explore the capabilities of current state-of-the-art LLMs on 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: and gain a deeper understanding of their strengths and limitations. To achieve this, we evaluate LLM performance across various aspects for each task. The details are shown in Appendix C.
For the FinSM task, we evaluate retrieval performance using Hit Rate (HR@kHR@k(He et al., 2017), Recall ($R@k$), and Macro-F1 ($MF1@k$), where k=1,5,10,20k={1,5,10,20}. Hit Rate measures whether the model retrieves at least one relevant item within the top-$k$ results. Recall assesses how completely the model retrieves all relevant elements. Macro-F1 balances precision and recall by averaging query-level F1 scores, providing an overall view of retrieval accuracy and coverage (Shani and Gunawardana, 2010).
For the FinRE task, we evaluate model performance using Accuracy (Acc), Macro Precision (P), Macro Recall (R), and Macro F1. Accuracy measures the overall proportion of correctly classified relations. Macro Precision and Macro Recall average performance across the three relation types. Macro F1 provides a balanced summary by combining precision and recall for each relation type and averaging the results.
For the FinMR task, we evaluated the mathematical reasoning capabilities of various LLMs using the LLM-as-a-judge approach (Gu et al., 2024; Li et al., 2024) to measure overall Accuracy. Additionally, we designed three error-rate indicators, including the structural error rate (SER), extraction error rate (EER), and calculation error rate (CER), to evaluate the quality of the generated answers by LLMs. Together, these metrics provide a comprehensive evaluation of mathematical reasoning performance.
4.4. Evaluation LLMs
Our goal is to evaluate the foundational capabilities of SOTA LLMs on the 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark, to better understand their strengths and limitations in financial auditing. To this end, we selected a diverse set of models: one closed-source general SOTA LLM (GPT-4o (Hurst et al., 2024)), ten open-source general SOTA LLMs of different sizes (DeepSeek-V3 (Liu et al., 2024), Qwen3-235B-A22B-Instruct-2507 (Team, 2025), Llama-4-Scout-17B-16E-Instruct (Meta, 2025), Qwen2.5-72B-Instruct (Yang et al., 2024), Llama-3.3-70B-Instruct (Grattafiori et al., 2024), Qwen3-32B (Team, 2025), gemma-3-27b-it (Team, 2024), gemma-3-12b-it (Team, 2024), Llama-3.1-8B-Instruct (Grattafiori et al., 2024), and Llama-3.2-3B-Instruct (Grattafiori et al., 2024)), and two open-source financial SOTA LLMs (Fino1 (Qian et al., 2025) and FinR1 (Liu et al., 2025)). Detailed specifications of the evaluated models, including the exact versions used, are provided in Table 9 in the Appendix.
4.5. Evaluation Environments
[…]
5. Results and Analysis
5.1. FinSM performance
Table 5 presents the Hit-Rate@k performance of various LLMs on the FinSM task under zero-shot settings. The results show that LLMs still struggle to detect semantically inconsistent US-GAAP tags, with the best model (DeepSeek-V3) achieving only 11.89% average hit rate. This highlights the challenge of aligning models with fine-grained financial semantics. Several open-source models, including DeepSeek-V3, Qwen3-235B, and gemma-3-27b-it, outperform the closed-source GPT-4o, indicating that open-source progress is closing the gap in domain-specific retrieval. In contrast, financial domain models (Fin-o1-14B and Fin-R1) perform poorly, with Fin-o1-14B failing to produce any correct hits. This suggests a mismatch between their generative training focus and the retrieval-oriented nature of FinSM. Qwen3-32B also yields zero hit rates, reflecting possible instability in mid-scale variants. Overall, model scale contributes to performance gains (e.g., Qwen3-235B vs. Qwen3-32B; gemma-27B vs. gemma-12B), but the Llama series shows inconsistent scaling, signifying the importance of architecture and alignment strategies.
Table 5. Hit-Rate@k (%) performance under the zero-shot settings on the FinSM task.
LLMs $HR@1$ $HR@5$ $HR@10$ $HR@20$ $HR@AVE$
GPT-4o 7.88 9.09 9.09 9.09 8.79
DeepSeek-V3 11.52 11.82 11.82 12.42 11.89
Qwen3-235B-A22B-Instruct-2507 9.09 10.00 10.00 10.00 9.77
Llama-4-Scout-17B-16E-Instruct 1.82 2.42 2.73 3.03 2.50
Qwen2.5-72B-Instruct 7.88 8.18 8.48 8.48 8.26
Llama-3.3-70B-Instruct 5.15 5.15 5.15 5.15 5.15
Qwen3-32B 0.00 0.00 0.00 0.00 0.00
gemma-3-27b-it 10.00 10.30 10.30 10.30 10.23
gemma-3-12b-it 8.79 9.70 9.70 10.30 9.62
Llama-3.1-8B-Instruct 6.06 6.06 6.06 6.06 6.06
Llama-3.2-3B-Instruct 3.33 3.94 4.24 4.55 4.02
Fin-o1-14B 0.00 0.00 0.00 0.00 0.00
Fin-R1 2.12 2.12 2.42 2.73 2.35
Table 6 reports the Recall@k performance of different LLMs on the FinSM task under zero-shot settings. Overall recall values remain low, with the best model, DeepSeek-V3, achieving only 8.83% on average. This suggests that while some models can identify relevant tags, they fail to comprehensively retrieve all mismatched US-GAAP elements within the top-$k$ predictions. Large open-source models such as DeepSeek-V3, Qwen3-235B, and gemma-3-27b-it outperform GPT-4o, showing growing competitiveness in capturing broader semantic coverage. In contrast, financial domain models (Fin-o1-14B and Fin-R1) perform poorly, with Fin-o1-14B yielding zero recall, indicating a mismatch between domain-specific training objectives and retrieval-oriented evaluation. Similarly, Qwen3-32B fails to retrieve relevant tags, reflecting inconsistency in mid-scale variants. As with the Hit Rate results, scaling trends are non-monotonic: larger models often perform better, but the Llama series exhibits irregular improvements. These results highlight the difficulty of FinSM from a recall perspective, which demands that models capture all semantically inconsistent tags across interrelated XBRL filings and taxonomy documents.
Table 6. Recall@k (%) performance under the zero-shot settings on the FinSM task.
LLMs $R@1$ $R@5$ $R@10$ $R@20$ $R@AVE$
GPT-4o 4.93 6.82 6.98 7.01 6.43
DeepSeek-V3 7.06 8.82 9.33 10.11 8.83
Qwen3-235B-A22B-Instruct-2507 6.38 7.77 8.17 8.33 7.66
Llama-4-Scout-17B-16E-Instruct 0.92 1.58 1.91 2.21 1.66
Qwen2.5-72B-Instruct 4.31 4.81 5.18 5.18 4.87
Llama-3.3-70B-Instruct 3.34 3.78 3.85 3.85 3.71
Qwen3-32B 0.00 0.00 0.00 0.00 0.00
gemma-3-27b-it 5.91 7.81 8.20 8.26 7.54
gemma-3-12b-it 5.09 7.06 7.61 8.49 7.06
Llama-3.1-8B-Instruct 2.88 4.32 4.73 4.82 4.19
Llama-3.2-3B-Instruct 1.64 3.23 3.68 3.98 3.13
Fin-o1-14B 0.00 0.00 0.00 0.00 0.00
Fin-R1 1.28 2.02 2.32 2.63 2.06
Table 7 presents the Macro-F1@k results for the FinSM task under zero-shot settings. Compared with Hit Rate and Recall, Macro-F1 provides a stricter measure by jointly considering precision and recall, reflecting the balance between correctly retrieved mismatched tags and false positives. Overall scores remain very low, with the best model reaching only 9.35% on average, underscoring the difficulty of achieving both accuracy and coverage in structured financial retrieval. DeepSeek-V3 performs best across all cutoffs, followed by Qwen3-235B and gemma-3-27b-it, which also outperform GPT-4o, showing that open-source models are increasingly competitive in balancing precision and recall. In contrast, financial domain models perform poorly, with Fin-o1-14B producing no correct outputs and Fin-R1 reaching only 1.84%, suggesting that domain adaptation without retrieval alignment may increase false positives. Scaling effects appear within some model families, but the Llama series shows inconsistent improvements, indicating that architecture and alignment strategies influence performance more than model size. These findings highlight that FinSM requires models not only to retrieve relevant tags but also to control spurious predictions and maintain precision-recall balance across structured, multi-document inputs.
Table 7. Macro-F1@k (%) performance under the zero-shot settings on the FinSM task.
LLMs $MF1@1$ $MF1@5$ $MF1@10$ $MF1@20$ $MF1@AVE$
GPT-4o 5.43 6.91 7.03 7.02 6.60
DeepSeek-V3 7.73 9.54 9.98 10.17 9.35
Qwen3-235B-A22B-Instruct-2507 6.85 7.83 7.96 7.97 7.65
Llama-4-Scout-17B-16E-Instruct 1.08 1.55 1.76 1.80 1.55
Qwen2.5-72B-Instruct 4.73 5.29 5.42 5.42 5.21
Llama-3.3-70B-Instruct 3.65 4.07 4.14 4.14 4.00
Qwen3-32B 0.00 0.00 0.00 0.00 0.00
gemma-3-27b-it 6.47 8.38 8.67 8.71 8.06
gemma-3-12b-it 5.57 7.62 8.04 8.27 7.38
Llama-3.1-8B-Instruct 3.29 4.47 4.77 4.80 4.33
Llama-3.2-3B-Instruct 2.00 2.88 2.95 2.94 2.69
Fin-o1-14B 0.00 0.00 0.00 0.00 0.00
Fin-R1 1.46 1.92 1.97 2.00 1.84
Overall, the results confirm that FinSM presents a highly challenging setting based on structured, multi-document inputs. Across Hit Rate, Recall, and Macro-F1, current LLMs struggle with both coverage and the balance between precision and recall. Large open-source models such as DeepSeek-V3, Qwen3-235B, and gemma-3-27b-it outperform GPT-4o, whereas financial domain models fail to adapt, showing that domain specialization alone does not guarantee retrieval effectiveness. FinSM therefore serves as a rigorous benchmark for evaluating LLMs’ capacity for semantic retrieval and taxonomy alignment in realistic financial auditing contexts.
5.2. FinRE performance
Table 8 reports the overall performance of various LLMs on the FinRE task under zero-shot settings. The results demonstrate that relationship extraction is highly challenging for current LLMs, as it requires precise interpretation of hierarchical and compositional dependencies within structured XBRL filings and their alignment with the US-GAAP taxonomy. Among all models, GPT-4o achieves the best performance with 91.82% accuracy and balanced precision, recall, and F1 (around 90%), showing a strong capability to consistently identify different types of erroneous relationships. DeepSeek-V3 follows with 82.73% accuracy and 80.68% macro-F1, which, while competitive, indicates a noticeable gap compared to GPT-4o in balancing precision and recall.
Table 8. The overall performance (%) under the zero-shot settings on the FinRE task.
LLMs Acc Macro P Macro R Macro F1
GPT-4o 91.82 90.15 90.03 90.09
DeepSeek-V3 82.73 87.81 80.19 80.68
Qwen3-235B-A22B-Instruct-2507 62.73 69.22 67.18 63.56
Llama-4-Scout-17B-16E-Instruct 27.50 17.09 35.82 22.71
Qwen2.5-72B-Instruct 67.50 72.14 71.50 68.27
Llama-3.3-70B-Instruct 28.86 23.06 38.26 21.54
Qwen3-32B 0.00 0.00 0.00 0.00
gemma-3-27b-it 45.91 35.97 53.00 40.85
gemma-3-12b-it 27.95 31.78 25.81 27.37
Llama-3.1-8B-Instruct 30.45 15.25 26.58 19.34
Llama-3.2-3B-Instruct 17.73 22.31 12.74 15.70
Fin-o1-14B 0.00 0.00 0.00 0.00
Fin-R1 32.73 34.05 21.00 25.97
By contrast, the majority of open-source models lag significantly. Qwen2.5-72B and Qwen3-235B achieve moderate performance (Acc 67.50% and 62.73%, respectively), suggesting partial capability in capturing structural errors but with reduced robustness. Mid-scale models such as gemma-3-27b-it (Acc 45.91%) and gemma-3-12b-it (Acc 27.95%) perform weakly, while the entire Llama series struggles across scales. In particular, Llama-4-Scout-17B-16E-Instruct, despite being an extremely large-scale model, reaches only 27.50% accuracy and 22.71% macro-F1, performing worse than some mid-scale Qwen and Gemma models. This highlights that scale alone does not translate into structural reasoning ability without proper alignment to taxonomy-driven tasks. Notably, Qwen3-32B and the financial domain-specific Fin-o1-14B completely fail (0.00% across all metrics), while Fin-R1 remains weak with only 32.73% accuracy and 25.97% macro-F1, underscoring that domain specialization without explicit structural training does not generalize to relationship extraction.
Figure 3 reveals substantial variation across relation types. GPT-4o demonstrates the most consistent performance, exceeding 80% F1 on Reversal and Inappropriateness and achieving 100% on CombinationErr, showing strong robustness across all categories. DeepSeek-V3 follows with balanced results, though consistently lower than GPT-4o. In contrast, most open-source models struggle on CombinationErr, with many collapsing to near zero, indicating that axis-member validation across linkbases poses the greatest challenge. The Llama family performs poorly across all types, and both Qwen3-32B and Fin-o1-14B completely fail, achieving 0.00 in every category. Interestingly, Fin-R1 reaches a moderate 60.51 on CombinationErr but remains weak elsewhere. These results suggest that while top-tier general models can handle structural relationship extraction, most open-source and financial-domain models lack the structural alignment needed, with CombinationErr emerging as the hardest category overall.
Figure 3. The F1-score (%) for individual relation type under the zero-shot settings on the FinRE task.
In summary, the FinRE results show that while frontier models such as GPT-4o achieve strong and balanced performance across all relation types, most open-source models lag considerably, and domain-specific financial LLMs fail to generalize to this structured classification setting. At the label level, CombinationErr emerges as the most difficult category, where even competitive models face a sharp drop, indicating that validating axis–member consistency requires reasoning across multiple interrelated documents. Overall, FinRE highlights that accurate financial relationship extraction demands more than scale or domain specialization, instead requiring explicit alignment with taxonomy-driven, cross-document reasoning.
5.3. FinMR performance
Figure 4. The accuracy (%) under the zero-shot settings on the FinMR task.
Figure 5. The error-rate results (%) for the FinMR task, where SER denotes the structural error rate, EER represents the extraction error rate, and CER refers to the calculation error rate.
Figure 4 reports the zero-shot accuracy on the FinMR task. Overall, all models show limited ability to perform structured mathematical reasoning, with the best model, Fin-o1-14B, reaching only 13.86% accuracy. This highlights the inherent difficulty of the task, which requires models to extract, compute, and verify numerical relations defined across hierarchical calculation structures in XBRL filings.
Across all LLMs, as shown in Figure 5, calculation errors (CER) dominate the performance gap. Most models exhibit extremely high CER values, typically between 70% and 83%, indicating that they struggle to follow multi-step arithmetic relationships or maintain consistency with domain-specific calculation rules. In contrast, extraction errors (EER) are comparatively moderate (roughly 10–22%), suggesting that LLMs can often locate the reported value but fail when performing reasoning-based computation.
Interestingly, Fin-o1-14B achieves the highest accuracy while maintaining a very low CER (9%), yet it suffers from a high structural error rate (SER = 71%). This pattern implies that once the model identifies a valid structural path, it performs reliable arithmetic reasoning but often fails to interpret the schema and linkbase structure correctly. Similarly, Qwen3-32B and Fin-R1 collapse due to severe structural errors, showing that without schema alignment, even accurate computation cannot be triggered.
Taken together, these results suggest that FinMR poses a fundamentally harder challenge than surface-level numerical QA. Success requires not only multi-step calculation ability but also schema-aware reasoning and structural grounding across multiple XBRL documents, capabilities that remain far beyond what current LLMs can reliably achieve.
6. Conclusion
In this work, we present 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first benchmark designed to evaluate LLMs on structured, hierarchical, and taxonomy-driven financial multi-documents. Built upon real-world XBRL filings, 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: defines three complementary tasks, FinSM, FinRE, and FinMR, which jointly assess structured semantic consistency, hierarchical relationship understanding, and multi-step mathematical reasoning. Our extensive zero-shot evaluation of SOTA LLMs shows that even leading models such as GPT-4o, DeepSeek-V3, and Fin-o1-14B still struggle with cross-document reasoning, structured interpretation, and taxonomy alignment. These findings highlight the limitations of current LLMs in financial auditing and emphasize the importance of developing models with stronger structural grounding and reasoning awareness. We release all datasets and evaluation code to support future research on trustworthy financial intelligence.
Appendix A LLM Judgment for Data Annotation
A.1. Prompt for the FinSM task
DQC_0099 Example You are given a DQC validation message. Your task is to extract two specific pieces of information: 1. **Statement Name**: Extract the name of the financial statement mentioned in the message. If it appears in a phrase like "0000007 - Statement - Consolidated Statements of Operations", ignore any prefixes such as "- Statement -" or ID numbers. Only return the clean name of the statement (e.g., "Consolidated Statements of Operations"). 2. **Erroneous us-gaap Tag(s)**: From all us-gaap:‘ prefixed tags mentioned in the message, extract only those that are clearly described as having an error or violation. A tag should be included **only if** it satisfies both of the following: - It is explicitly stated to be **not included in any calculation relationship**. - The message states that it **will produce an error**, **violates a rule**, or **should have been reported differently**. Do not include us-gaap:‘ tags that are merely suggested, referenced as valid abstract parents, or shown as examples for correction. 3. **Reasoning**: For each extracted erroneous tag, provide a short explanation summarizing why it was classified as erroneous, based on the message. Output your result as a structured JSON in the following format: ‘‘‘json { "statement_name": "...", "error_tags": [ { "tag": "...", "reason": "..." } ] } ‘‘‘
DQC_0109 Example You are given a DQC validation message. Your task is to extract all Dimensions information from the message. 1. Extract all key-value pairs under Dimensions, where each dimension is formatted as us-gaap:XXX=YYY. 2. For each extracted dimension, provide a short explanation in the "reason" field based on the message context. If no explicit issue is mentioned about a specific dimension, you may leave the "reason" as an empty string. Output your result as a structured JSON in the following format: ‘‘‘json { "error_tags": [ { "dimension": "...", "reason": "..." } ] } ‘‘‘
DQC_0137 Example You are given a DQC validation message. Your task is to extract the following information: 1. Target Tag: Extract the primary us-gaap tag being referenced or discussed as erroneous in the message. Only include the tag explicitly stated as problematic or requiring correction. 2. Dimensions: Extract all dimension key-value pairs listed under "Dimensions" in the message. Each dimension should be in the format us-gaap:XXX=YYY. 3. For each extracted dimension, provide a short explanation in the "reason" field based on the message context. If the message does not explicitly mention an issue with the dimension, you may leave the "reason" field empty or give a general summary. Output your result as a structured JSON in the following format: ‘‘‘json { "target_tag": "...", "error_tags": [ { "dimension": "...", "reason": "..." } ] } ‘‘‘
A.2. Prompt for the FinRE task
DQC_0081 Example You are given a DQC validation message. Your task is to extract two specific pieces of information: 1. **Statement Name**: Extract the name of the financial statement mentioned in the message. If it appears in a phrase like "00000002 - Statement - BALANCE SHEET", ignore any prefixes such as ID numbers or "- Statement -". Only return the clean name of the statement (e.g., "BALANCE SHEET"). 2. **Erroneous us-gaap Tag Relationships**: From all us-gaap:‘ tags mentioned in the message, identify pairs of tags where the relationship between them is described as incorrect or inconsistent. A tag pair should be extracted **only if** it satisfies both of the following conditions: - The two tags are stated to be in a specific structural relationship in the filers taxonomy (e.g., sibling, parent-child). - The message clearly indicates that this relationship contradicts the definition in the official US-GAAP taxonomy, or requires reviewer attention for potential misclassification. 3. **Reasoning**: For each extracted tag pair, provide a short explanation summarizing the inconsistency or incorrect relationship between them as described in the message. Output your result as a structured JSON in the following format: ‘‘‘json { "statement_name": "...", "error_tags": [ { "tag1": "...", "tag2": "...", "reason": "..." } ] } ‘‘‘
DQC_0001 Example You are an expert in XBRL taxonomies, tasked with parsing DQC validation messages for relationship errors (Rule DQC.US.0001), specifically an axis with an inappropriate member. 1. **Main Concept**: Extract the primary concept that is being dimensionally qualified. This is usually the first concept mentioned in the message. 2. **Dimension Pair**: From the Dimensions field in the message, extract the full axis and member combination string **exactly as it appears**, including all prefixes. 3. **Reasoning**: Provide a short explanation that the member is unallowable for the specified axis, as described in the message. Output your result as a single, structured JSON object in the following format: ‘‘‘json { "main_concept": "UnrealizedGainLossOnCashFlowHedgingInstruments", "dimension_pair": "us-gaap:FairValueByFairValueHierarchyLevelAxis=us-gaap:FairValueMeasurementsRecurringMember", "reason": "The message indicates the member is unallowable for the specified axis." } ‘‘‘
DQC_0145 Example You are an expert in XBRL taxonomies, tasked with parsing DQC validation messages for relationship errors (Rule DQC.US.0145), specifically Inappropriate Cash Flow Presentation’. 1. **Head Concept**: Extract the primary concept that is inappropriately presented. This is usually the first concept mentioned in the message. 2. **Tail Concept**: Extract the presentation concept that the head concept is incorrectly a descendant of. 3. **Reasoning**: Provide a short explanation that the head concept should not be presented as a component of the tail concept. Output your result as a single, structured JSON object in the following format: ‘‘‘json { "head_concept": "element 1", "tail_concept": "element 2", "reason": "The message indicates the head concept is inappropriately presented as a descendant of the tail concept and should be outside this group." } ‘‘‘
A.3. prompt for the FinMR task
DQC_0015 Example You are given a DQC validation message. Your task is to extract the following key information: 1. **Erroneous Tag Information:**: Identify the primary us-gaap tag in the message that is reported with an incorrect value due to sign (positive/negative) error. For this tag, extract: - tag‘: The us-gaap element tag with the error. - period‘: The reporting period for this tag, as shown in the "Period" field. - reported_value‘: The actual value reported in the filing, as stated in the message or "Value" field. - correct_value‘: The correct value, which should be the absolute value of the reported value (i.e., reported_value without the negative sign), as implied or explicitly required by the message. - reason‘: Briefly explain why the reported value is considered incorrect, based on the message content (for example, "This element should not have a negative value. The amount should be input as a positive value and, if necessary, a negated label should be provided."). Return your result as structured JSON in the following format: ‘‘‘json { "error_tags": { "tag": "...", "period": "...", "reported_value": "...", "correct_value": "...", "reason": "..." } } ‘‘‘
DQC_0117 Example You are an expert financial analyst parsing DQC validation messages for mathematical reasoning errors (Rule DQC.US.0117). Your task is to extract all required components from the message. 1. **Statement Name**: Extract the clean name of the financial statement (e.g., "Condensed Consolidated Income Statements"). 2. **Input Data**: From the properties list at the end of the message, extract the following: * target_concept‘: The full name of the primary us-gaap:‘ concept being evaluated. * period‘: The reporting period for the fact (e.g., "2020-10-01 to 2020-12-31"). 3. **Output Data**: From the main body of the message, extract the following values: * extracted_value‘: The value of the concept as reported in the filing (e.g., "255,500,000"). * calculated_value‘: The correct value based on the dimensional breakdown (e.g., "142,400,000"). * is_correct‘: This should always be "No" for these error messages. Output your result as a single, structured JSON object in the following format: ‘‘‘json { "statement_name": "...", "input": { "target_concept": "us-gaap:RevenueFromContractWithCustomerIncludingAssessedTax", "period": "..." }, "output": { "extracted_value": "...", "calculated_value": "...", "is_correct": "No" } } ‘‘‘
DQC_0126 Example You are given a DQC validation message. Your task is to extract the following key information: 1. **Statement Name**: Extract the name of the financial statement mentioned in the message. If it appears in a phrase like "000004 - Statement - CONSOLIDATED STATEMENTS OF OPERATIONS (Unaudited)", ignore any numeric prefixes and the phrase "- Statement -". Only return the clean name of the statement, such as "CONSOLIDATED STATEMENTS OF OPERATIONS (Unaudited)". 2. **Erroneous Total Element Tag Information**: Extract detailed information about the total element (‘us-gaap:‘ tag) that is reported incorrectly. Specifically, you should extract: - tag‘: The total element tag, as indicated by the **"Total Element"** field in the message. - period‘: The reporting period associated with this tag, as indicated by the **"Total period"** field. - reported_value‘: The value that was reported in the filing, taken from the **"Total Value"** field. - correct_value‘: The expected value, i.e., the **sum of the child components** as defined in the calculation linkbase. - This is usually explicitly stated in a sentence like: *"The sum of these child components is..."* - You may cross-validate this by summing the listed child component values in the message if provided. - reason‘: Provide a concise explanation of the error - why the reported value is considered incorrect - based directly on the message content. Only include this tag if all of the following are true: - The message identifies a total element with incorrect calculation - Both reported and correct values are stated or inferable - The error occurs in a specified reporting period Return your result as structured JSON in the following format: ‘‘‘json { "statement_name": "...", "error_tags": { "tag": "...", "period": "...", "reported_value": "...", "correct_value": "...", "reason": "..." } } ‘‘‘
Appendix B US-GAAP Taxonomy Chunking
To enable structured retrieval and reasoning, we convert the official US-GAAP taxonomy spreadsheets into concept-centric chunks. Each GAAP concept is represented by a core chunk capturing its intrinsic attributes and a set of relation chunks describing its semantic and structural links to other concepts. The conversion operates directly on the official Excel workbook through systematic normalization and extraction.
Sheet Normalization.
All sheets are loaded from the taxonomy workbook, column headers are standardized, and relevant sheets (Concept, Presentation, Calculation, Definition, Reference, Enumeration) are identified by keyword matching. Unnamed or irrelevant reference sheets are excluded, and each extracted element is annotated with provenance information such as file name, sheet name, and row number.
Core Chunk Extraction.
From the Concepts sheet, each concept’s identifier, label, type, balance, period type, abstract flag, documentation, and deprecation status are extracted and rendered into a canonical text form (concept::core) suitable for retrieval or embedding.
Relation Chunk Extraction.
For each concept, all relation-bearing sheets are parsed to construct chunks representing presentation hierarchies, calculation formulas, definition arcs, reference citations, and enumerations. The parser handles heterogeneous layouts, missing columns, and varying role conventions. Each chunk is recorded with its role, arcrole, and provenance metadata under the identifier concept::relations:pres—calc—def—ref—enum, with integrity checks ensuring that all relation targets correspond to valid core concepts.
Outputs.
All chunks are written into chunks_core.jsonl and chunks_relations.jsonl, and summary statistics are stored in meta.json. This representation preserves the hierarchical and relational structure of the US-GAAP taxonomy while making each concept independently retrievable.
To enable structured retrieval and reasoning, we convert the official US-GAAP taxonomy spreadsheets into concept-centric chunks. Each GAAP concept is represented by two complementary components: a core chunk that captures intrinsic attributes and a collection of relation chunks that encode its semantic and structural links to other concepts. The conversion process operates directly on the official Excel workbook, performing systematic normalization and extraction as follows.
Sheet Normalization and Identification.
We first load all sheets within the taxonomy workbook, standardize their column headers by lowercasing and trimming whitespace, and identify relevant sheets by keyword matching (Concept, Presentation, Calculation, Definition, Reference, and Enumeration). Suspicious “all-unnamed” reference sheets are ignored. The taxonomy version number is automatically inferred from the filename. Each extracted element is annotated with its provenance, including the file name, sheet name, and row number.
Core Chunk Extraction.
From the Concepts sheet, we extract for each concept its identifier (prefix:name), label, type, balance, period type, abstract flag, documentation, and deprecation status. These attributes are rendered into a canonical, human-readable chunk_text and stored with a stable identifier in the form concept::core. The core chunk provides a self-contained textual summary suitable for direct retrieval or embedding.
Relation Chunk Extraction.
For every concept, we parse all relation-bearing sheets to construct one chunk per relation family:
  • Presentation: hierarchical parent–child paths with associated roles and preferred labels.
  • Calculation: quantitative relationships including parent/child direction, weights, and roles.
  • Definition: arcs between dimensions, domains, and hypercubes, including arcrole, source, and target.
  • Reference: external standard citations consisting of source, section, and note, filtered to the relevant concept.
  • Enumeration: extensible lists aggregated by domain and linkrole, with de-duplicated members.
The extraction is designed to handle heterogeneous sheet layouts and missing columns: when explicit from*/to* fields are absent, the parser falls back to parent/child-style columns and merges multiple role or linkrole conventions. Each relation chunk records the roles and arcroles encountered, as well as complete provenance metadata, and uses the identifier concept::relations:{pres—calc—def—ref—enum} to store. Integrity validation ensures that all relation targets reference existing core concepts, and malformed rows are automatically skipped.
Outputs.
All chunks are written into chunks_core.jsonl and chunks_relations.jsonl, while global statistics are summarized in meta.json. This chunked representation preserves the hierarchical and relational organization of the US-GAAP taxonomy while making each concept addressable as an atomic, retrievable unit, as shown in the following.
Example of a Core Chunk [Concept Core] ID: us-gaap:Revenues Label: Revenues Type: monetaryItemType | Balance: credit | PeriodType: duration | Abstract: False Status: active | DeprecatedLabel: None | DeprecatedDate: None Documentation: Amount of revenue recognized from goods sold or services rendered during the period. Provenance: file=GAAP_Taxonomy_2024.xlsx | sheet=Concepts | row=248
Example of a Relation Chunk [Concept Relations] ID: us-gaap:Revenues Presentation: - Role: Statement - Income Statement Path: Revenues > CostOfRevenue > GrossProfit Calculation: - As Parent: child=us-gaap:CostOfRevenue | weight=-1.0 | role=Calculation Linkbase - As Parent: child=us-gaap:GrossProfit | weight=+1.0 | role=Calculation Linkbase Definition (Dimensions/Domain/Hypercube): - domain-member: from=RevenueRecognitionPolicyTextBlock -> to=RevenueCategoryAxis References: - FASB ASC 606 | Section 25 | Revenue Recognition Provenance: sheets include Presentation-like and Calculation-like sheets
Appendix C Evaluation Metrics
C.1. The metrics for FinSM
For the FinSM task, we evaluated retrieval performance using Hit Rate ($HR@k$), Recall ($R@k$), and Macro-F1 ($MF1@k$), where $k=\{1,5,10,20\}$. Hit Rate evaluates whether the LLM retrieves at least one relevant element within the top-$k$ results. It reflects the proportion of queries for which the model can successfully hit a relevant item, without considering how many are retrieved. Recall measures the fraction of all relevant elements that are successfully retrieved within the top-$k$ predictions. It captures how completely the model covers the relevant set under a cutoff of $k$. Macro-F1 balances precision and recall at the query level by computing the F1 score for each query using its top-$k$ predictions and then averaging across all queries. It reflects both the accuracy (precision) and completeness (recall) of retrieval, giving equal weight to each query.
(5) $\scriptsize\mathrm{HR}@k=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\big(E_{i}^{(k)}\cap G_{i}\neq\emptyset\big)$
(6) $\scriptsize\mathrm{R}@k=\frac{1}{N}\sum_{i=1}^{N}\frac{|E_{i}^{(k)}\cap G_{i}|}{|G_{i}|}$
(7) $\scriptsize\mathrm{Macro\text{-}F1}@k=\frac{1}{N^{\prime}}\sum_{i=1}^{N^{\prime}}\frac{2\times P_{i}^{(k)}\times R_{i}^{(k)}}{P_{i}^{(k)}+R_{i}^{(k)}}$
where
$P_{i}^{(k)}$ $=\begin{cases}\dfrac{|E_{i}^{(k)}\cap G_{i}|}{|E_{i}^{(k)}|},&|E_{i}^{(k)}|>0\\ 0,&\text{otherwise}\end{cases},\qquad R_{i}^{(k)}=\begin{cases}\dfrac{|E_{i}^{(k)}\cap G_{i}|}{|G_{i}|},&|G_{i}|>0\\ 0,&\text{otherwise}\end{cases}$
and $N$ is the total number of queries, and $N^{\prime}$ is the number of queries with non-empty ground truth sets. $E_{i}^{(k)}$ denotes the top-$k$ predicted elements for the $i$-th query. $G_{i}$ denotes the set of ground truth elements for the $i$-th query. $|E_{i}^{(k)}\cap G_{i}|$ is the number of correctly predicted elements among the top-$k$ results. $k$ is the cutoff rank indicating the number of top predictions considered. $\mathbf{1}(\cdot)$ is the indicator function, returning 1 if the condition holds and 0 otherwise.
C.2. The metrics for FinRE
For the FinRE task, we assessed the performance of various LLMs using Accuracy (Acc), Precision (Macro P), Recall (Macro R), and F1-score (Macro F1). Accuracy measures the overall proportion of correctly classified relations among all predictions. Precision evaluates the proportion of correctly predicted relations within each class, averaged across the three relation categories (Macro P). Recall quantifies the proportion of true relations that were successfully identified for each class, also averaged across classes (Macro R). F1-score provides a balanced measure of Precision and Recall by computing their harmonic mean per class and then averaging across the three classes (Macro F1).
(8) $\scriptsize\mathrm{Acc}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}(\hat{y}_{i}=y_{i})$
(9) $\scriptsize\mathrm{Macro\text{-}P}=\frac{1}{C}\sum_{c=1}^{C}\frac{TP_{c}}{TP_{c}+FP_{c}}$
(10) $\scriptsize\mathrm{Macro\text{-}R}=\frac{1}{C}\sum_{c=1}^{C}\frac{TP_{c}}{TP_{c}+FN_{c}}$
(11) $\scriptsize\mathrm{Macro\text{-}F1}=\frac{1}{C}\sum_{c=1}^{C}\frac{2\times P_{c}\times R_{c}}{P_{c}+R_{c}}$
where $N$ is the total number of instances. $C$ is the number of relation categories (here $C=3$). $y_{i}$ and $\hat{y}_{i}$ denote the ground truth and predicted relation labels for the $i$-th instance. $TP_{c}$, $FP_{c}$, and $FN_{c}$ denote the number of true positives, false positives, and false negatives for class $c$, respectively. $P_{c}$ and $R_{c}$ denote the precision and recall for class $c$. $\mathbf{1}(\cdot)$ is the indicator function that equals 1 when the prediction is correct and 0 otherwise.
C.3. The metrics for FinMR
For the FinMR task, we evaluated the mathematical reasoning capabilities of various LLMs using an LLM-as-a-judge framework to assess the overall Accuracy of generated answers. This metric reflects whether each model produces the correct numerical result under the specified auditing rule. In addition to overall correctness, we introduced three fine-grained error indicators, namely the Structural Error Rate (SER), Extraction Error Rate (EER), and Calculation Error Rate (CER), each corresponding to a specific stage of reasoning validation.
SER measures the proportion of outputs with invalid structure, where the model fails to produce the required JSON format containing both the extracted and calculated values. EER quantifies errors in identifying or reproducing the correct extracted value, meaning that the predicted numerical entity does not match the true extracted value in mathematical meaning. CER captures computational mistakes, where the predicted calculated value deviates from the ground truth even when the structure and extracted value are both correct. Together, these indicators provide a hierarchical view of reasoning quality, reflecting whether an error arises from structural formatting, factual extraction, or numerical computation, thereby offering fine-grained insights into model performance on complex financial mathematical reasoning tasks. The specific prompt for judgment is shown below.
LLM-as-a-judge prompt for FinMR Task Instruction: You are an evaluator. Your task is to judge whether the models output pred_answer is correct compared to the given true_answer. Follow the rules strictly: Step 1 (Structure Check): Verify whether pred_answer has the same structure as true_answer. The required structure is a JSON object with exactly two keys: {{"extracted_value": <value>, "calculated_value": <value>}} Minor formatting differences (e.g., line breaks, indentation, whitespace) are acceptable. If the structure is invalid, output the label: S If valid, continue to Step 2 Step 2 (Extracted Value Check): Compare true_answer["extracted_value"] and pred_answer["extracted_value"] by their mathematical meaning, not their string form. For example, "-1,284" and "-1284" are considered equal. If they are not equal in numeric meaning, output the label: E If equal, continue to Step 3 Step 3 (Calculated Value Check): Compare true_answer["calculated_value"] and pred_answer["calculated_value"] strictly in numeric meaning. They must be exactly equal (zero tolerance). If they are not equal, output the label: C If equal, then everything is correct Final Decision: If all three checks pass, output the label: A Output only one label: S, E, C, or A. Do not explain your reasoning.
Given a total of $N$ evaluated instances, let $N_{A}$, $N_{S}$, $N_{E}$, and $N_{C}$ denote the numbers of instances labeled as Accurate (A), Structural Error (S), Extraction Error (E), and Calculation Error (C), respectively. Let $N_{\text{valid}}$ represent the number of successfully parsed instances that received a valid label. We define the following metrics:
(12) $\scriptsize\text{Accuracy (ACC)}=\frac{N_{A}}{N_{\text{valid}}}$
(13) $\scriptsize\text{Structural Error Rate (SER)}=\frac{N_{S}}{N_{\text{valid}}}$
(14) $\scriptsize\text{Extraction Error Rate (EER)}=\frac{N_{E}}{N_{\text{valid}}}$
(15) $\scriptsize\text{Calculation Error Rate (CER)}=\frac{N_{C}}{N_{\text{valid}}}$
where each instance is judged by the LLM-as-a-judge framework to output exactly one label from $\{A,S,E,C\}$. ACC, SER, EER, and CER, respectively, quantify the proportions of correct, structural, extraction, and calculation outcomes among validly parsed cases.
Appendix D Evaluation Models
Table 9 summarizes all models evaluated in this study, categorized by openness, domain specialization, and architectural foundation. The evaluation covers a wide range of both closed- and open-source large language models (LLMs) as well as domain-specific and pretrained baselines.
Table 9. Model categories and corresponding repositories.
Model Size Repository
Closed-source LLMs
GPT-4o gpt-4o-2024-08-06
Open-source LLMs
DeepSeek-V3 685B deepseek-chat
Llama-4-Scout-17B-16E-Instruct 109B meta-llama/Llama-4-Scout-17B-16E-Instruct
Qwen3-235B-A22B-Instruct-2507 235B Qwen/Qwen3-235B-A22B-Instruct-2507
Llama-3.3-70B-Instruct 70B meta-llama/Llama-3.3-70B-Instruct
Qwen2.5-72B-Instruct 72B Qwen/Qwen2.5-72B-Instruct
Qwen3-32B 32B Qwen/Qwen3-32B
gemma-3-27b-it 27B google/gemma-3-27b-it
gemma-3-12b-it 12B google/gemma-3-12b-it
Llama-3.1-8B-Instruct 8B meta-llama/Llama-3.1-8B-Instruct
Llama-3.2-3B-Instruct 3B meta-llama/Llama-3.2-3B-Instruct
Financial-specific LLMs
Fin-o1-14B 14B TheFinAI/Fin-o1-14B
Fin-R1 7B SUFE-AIFLM-Lab/Fin-R1
  • Closed-source LLMs: We include GPT-4o (Hurst et al., 2024), accessed via OpenAI’s API, as a representative of state-of-the-art proprietary models. Although the architecture and model size remain undisclosed, GPT-4o serves as a strong upper-bound reference in our benchmark.
  • Open-source LLMs: This category encompasses a diverse set of publicly available, instruction-tuned models, including DeepSeek-V3 (Liu et al., 2024), Qwen3-235B-A22B-Instruct-2507 (Team, 2025), Qwen2.5-72B-Instruct (Yang et al., 2024), and Qwen3-32B (Team, 2025), offering a wide scale range. We also include several Llama models from Meta: Llama-4-Scout-17B-16E-Instruct (Meta, 2025), Llama-3.3-70B-Instruct (Grattafiori et al., 2024), Llama-3.1-8B-Instruct (Grattafiori et al., 2024), and Llama-3.2-3B-Instruct (Grattafiori et al., 2024), as well as Google’s Gemma models (Team, 2024) (gemma-3-27b-it and gemma-3-12b-it). These models collectively enable cross-architecture and scaling comparisons across modern open-source systems.
  • Financial-specific LLMs: We include Fin-o1-14B (Qian et al., 2025) and Fin-R1 (Liu et al., 2025), both trained on domain-specific financial corpora to capture specialized terminology and reasoning patterns in financial reporting. These models allow us to assess the impact of domain adaptation on structured auditing and numerical reasoning tasks.
Together, these models form a comprehensive evaluation spectrum, spanning closed- and open-source systems, general-purpose and domain-specialized models, and encoder- versus decoder-based architectures, enabling an in-depth analysis of performance across all proposed benchmark tasks.